OpenStudy (anonymous):

solve (3a-4b)/2 = c for letter b

OpenStudy (anonymous):

3a-4b=2c 4b=3a-2c b= 1/4(3a-2c)

OpenStudy (anonymous):

3a-4b=2c 4b=3a-2c Multiply throughout by (-) then divide throughout by 4 b= 1/4(3a-2c)